# When higher-order interactions enhance synchronization: the case of the Kuramoto model

### v1

@@ Acknowledgments
Figure SM4: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
@@
Figure SM5: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
@@
Figure SM6: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).

### v2

@@ Acknowledgments
Figure SM4: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
@@
Figure SM5: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
@@
Figure SM6: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM7: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ distributed normally with average $0.15$ and standard deviation $0.1$, and initial phases distributed uniformly in $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
Figure SM8: System response to different pairwise ($K_{1}$) and higher-order ($K_{2}$) coupling strengths, for initially incoherent $1$-connected hypergraphs from Fig. SM7. The top-center panel replicates the bottom right panel of Fig. SM7 for comparison (color is the mean order parameter $R$). The top-left (resp. top-right) panel shows the mean value of $R$ (blue line) and its standard deviation (shaded area) as a function of the higher-order coupling strength $K_{2}$ (resp. $K_{1}$) while keeping the pairwise interaction strength $K_{1}$ (resp. $K_{2}$) fixed. The values of $K_{1}$, $K_{2}$ explored are also indicated in the top-center panel by a horizontal (resp. vertical) black line. Pink (resp. brown) triangles pointing upward/downward (resp. leftward/rightward) mark representative pairs $(K_{1},K_{2})$. The bottom panels display the distributions (integral normalized to 11) of $R$ across the $N_{\mathcal{H}}=300$ realizations of hypergraph, initial conditions, and frequencies.
Figure SM9: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ distributed normally with average $0.15$ and standard deviation $0.1$, and initial phases distributed uniformly in $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
Figure SM10: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ distributed normally with average $0.15$ and standard deviation $0.1$, and initial phases distributed uniformly in $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R_{2}$.
Figure SM11: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ distributed normally with average $0.15$ and standard deviation $0.1$, and initial phases distributed uniformly in $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R_{2}$.
Figure SM12: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies distributed normally with average $0.15$ and standard deviation $0.1$ and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM13: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies distributed normally with average $0.15$ and standard deviation $0.1$ and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM14: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies distributed normally with average $0.15$ and standard deviation $0.1$ and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM15: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies distributed normally with average $0.15$ and standard deviation $0.1$ and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM16: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=100$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $4\times 7$. Each hypergraph has $N_{0}=100$ nodes, $N_{1}=2200$ links and $N_{2}=13{,}475$ triangles. On the left panel, oscillator frequencies $\omega_{j}$ are uniformly distributed in $[0,0.3]$, while on the right panel they are distributed normally with mean $0.15$ and standard deviation $0.1$. Initial phases are drawn uniformly in $0,\theta_{0}^{\mathrm{max}}=2\pi]$.
Figure SM17: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=100$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $4\times 7$. Each hypergraph has $N_{0}=100$ nodes, $N_{1}=2200$ links and $N_{2}=13{,}475$ triangles. On the left panel, oscillator frequencies $\omega_{j}$ are uniformly distributed in $[0,0.3]$, while on the right panel they are distributed normally with mean $0.15$ and standard deviation $0.1$. Initial phases are drawn uniformly in $0,\theta_{0}^{\mathrm{max}}=2\pi]$.
Figure SM18: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $31\times 31$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=100$ random hypergraphs were generated, with $N_{0}=100$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panel, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panel. The relative link cost is fixed at $c_{1}=1$, while triangle cost is $c_{2}=3$. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM19: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $31\times 31$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=100$ random hypergraphs were generated, with $N_{0}=100$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panel, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panel. The relative link cost is fixed at $c_{1}=1$, while triangle cost is $c_{2}=3$. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM20: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $31\times 31$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=100$ random hypergraphs were generated, with $N_{0}=100$ nodes, frequencies distributed normally with mean $0.15$ and standard deviation $0.1$, and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panel, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panel. The relative link cost is fixed at $c_{1}=1$, while triangle cost is $c_{2}=3$. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM21: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $31\times 31$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=100$ random hypergraphs were generated, with $N_{0}=100$ nodes, frequencies distributed normally with mean $0.15$ and standard deviation $0.1$, and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panel, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panel. The relative link cost is fixed at $c_{1}=1$, while triangle cost is $c_{2}=3$. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).